PLease help me write v in the form ai + bj + ck and find ||V|| if the initial point p is (4,7,0) and the terminal point Q is (0,5,6).
Thank you, Ray

Answer by Alan3354(8354)

PLease help me write v in the form ai + bj + ck and find ||V|| if the initial point p is (4,7,0) and the terminal point Q is (0,5,6).
==============
a = 0 - 4 (i of Q - P)
b = 5 - 7
c = 6 - 0
So V = -4i - 2j + 6k
---------------
The ABS of V is sqrt(a^2 + b^2 + c^2)
= sqrt(16 + 4 + 36)
= sqrt(56)
||V|| = 2*sqrt(14)